Gelbart Company manufactures gas grills. Fixed costs amount to $27,460,800 per year. Variable costs per gas grill are $480, and the average price per gas grill is $1,200.

Required: 1. How many gas grills must Gelbart Company sell to break even?

??? gas grills

2. If Gelbart Company sells 41,830 gas grills in a year, what is the operating income?

$ 3. If Gelbart Company’s variable costs increase to $504 per grill while the price and fixed costs remain unchanged, what is the new break-even point? If required, round your answer to the nearest whole number.

??? gas grills
Answer #1
Required 1 :
Contribution margin per unit = Selling price per unit - Variable cost per unit = 1200 - 480    720
Break even point = Fixed costs / Contribution margin per unit = 27460800 / 720 38140 gas grills
Required 2 :
Sales ( 41830 * 1200 ) 50196000
(-) Variable cost ( 41830 * 480 ) 20078400
Contribution margin 30117600
(-) Fixed cost 27460800
Operating income 2656800
Required 3 :
Contribution margin per unit = Sales price per unit - Variable cost per unit = 1200 - 504    696
Break even point = Fixed costs / Contribution margin per unit = 27460800 / 696 39455 gas grills